OpenStudy (anonymous):

Algebra 2 essay questions? someone please help?

OpenStudy (anonymous):

Show all work in simplifying \[\frac{ 3x^2-3x-60 }{ x^2+9x+20 }\] Be sure to list restrictions. Use complete sentences to explain your process.

OpenStudy (anonymous):

\[3-\frac{ 30 }{ x+5 }\]

OpenStudy (anonymous):

\[\frac{ 3(x-5) }{ x+5 }\]

OpenStudy (anonymous):

\[\frac{ 3x-15 }{ x+5 }\]

OpenStudy (anonymous):

?? :(

OpenStudy (tkhunny):

It seems like you are doing things by magic. Consider the Numerator by itself. Factor it and show the final result. Consider the Denominator by itself. Factor it and show the final result.

OpenStudy (anonymous):

x=5?

OpenStudy (anonymous):

I have no idea how to do this whole process :(

OpenStudy (tkhunny):

It is not clear what you are doing or trying to do. Do FIRST, the two steps I suggested, above. Factor, factor, factor and show your work. Don't try to do ANYTHING else. Just factor the Numerator and Denominator and provide your result.

OpenStudy (anonymous):

But I don't know how :( I missed a few lessons

OpenStudy (tkhunny):

It looks to me like you have a clue. \(3x^{2} - 3x - 60\;=\;3(x^{2} - x - 20)\;=\;3(x-5)(x+4)\) Can you do that? Demonstrate the denominator.

OpenStudy (anonymous):

That's how I did it. I needed to look at what I would come up with. I don't know how to get there from here though.

OpenStudy (tkhunny):

That isn't "did it" at all. Your task is to simplify the expression. Thre is no "y =". There is no graph. What you need to do is FACTOR. Do you see what I did in the Numerator? I started with the original expression and FACTORED. Please forget the graph from Wolfram. Please forget everything else but factoring the Denominator. Go!

OpenStudy (anonymous):

But I do not know how to factor the denominator.

OpenStudy (tkhunny):

Well, that is a problem, since there is no other effective method to simplify this expression. If you have NO background in factoring quadratic trinomials, you cannot solve this problem that requires you to simplify the expression. That being said, why do you have no idea how to factor a quadratic trinomial? You should have worked at least 50 of these before you ever were presented with this problem. Thus, where did this one break down? Have you been attending class? Did you do TOO well on a placement exam and get into the wrong class? \( x^{2} + 9x + 20 = (x + What)(x + What)\) Fill in the "What"s.